OpenStudy (anonymous):

determine the limit of : (3n+4)(1-n)/n^2

OpenStudy (anonymous):

when n goes to what

OpenStudy (anonymous):

n is raised to the 2nd power, n to infinity

OpenStudy (amistre64):

expand the top

OpenStudy (anonymous):

-3

OpenStudy (amistre64):

the only important terms are 3n*(-n) and n^2

OpenStudy (amistre64):

-3n^2/n^2 = -3 for all sufficiently large values of 3 :)

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}\frac{(3n+4)(1-n)}{n^2}\] \[\lim_{n \rightarrow \infty}\frac{3n(-n)+\cdots blah}{n^2}\]

OpenStudy (anonymous):

-3n^2-n+4/n^2 n^2(-3n-1/n+4/n^2)/n^2=-3n-1/n+4/n^2 when n goes to infinity -1/n+4/n^2goes to 0, only thing left is -3n lim as n goes to infinity -3n=-infinity

OpenStudy (anonymous):

gracias